When the son will be as old as father is today,their ages will add upto126 years.If the present age of father is 52 years. Find the present age of son.
Answer by ankor@dixie-net.com(22740)

When the son will be as old as father is today, their ages will add up to 126 years.
If the present age of father is 52 years, Find the present age of son.
:
52 + f = 126
f = 126 - 52
f = 74 yrs is the age of the father when son is 52
The difference in their ages 74 - 52 = 22 yrs
then
s = 52 - 22
s = 30 yrs is the son's present age